OpenStudy (anonymous):

What is the value of the expression 7^–3?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

7^-3 is the problem

OpenStudy (anonymous):

\[7^{-3} = \frac{ 1 }{ 7 } *\frac{ 1 }{ 7 }*\frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

1/343?

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

Correct. :)

OpenStudy (anonymous):

yay more?

OpenStudy (anonymous):

sure. :)

OpenStudy (anonymous):

What is the value of the expression

OpenStudy (anonymous):

\[\frac{ 1 }{4^{-3}} = \frac{ 1 }{ \frac{ 1 }{ 4^3 } } = 4^3\]

OpenStudy (anonymous):

answer choices:A. 12 B. 64 C.1/64 D.1/12

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

4*4* 4 = ?

OpenStudy (anonymous):

64